OpenStudy (anonymous):

MEDAL REWARDED Find the zeros of the function. f(x) = 9x3 + 36x2 + 27x

OpenStudy (anonymous):

@John_ES

hero (hero):

When it says to find the "zeroes of the function", it wants you to set \(f(x) = 0\), then solve for \(x\). In this case, \(0 = 9x^3 + 36x^2 + 27x\)

hero (hero):

You can start off by factoring \(x\) from each term on the right side. \(0 = x(9x^2 + 36x + 27)\)

hero (hero):

Now you have a quadratic in the parentheses that you can factor

hero (hero):

Actually, we can factor out 9 as well to get: \(0 = 9x(x^2 + 4x + 3)\)

hero (hero):

That should make the quadratic expression inside the parentheses easier to factor.

OpenStudy (anonymous):

And then what? sorry i learn better when someone just does the whole thing and then i look at their steps

hero (hero):

So next we have to factor the quadratic expression inside the parentheses: \(x^2 + 4x + 3\) Can you factor that or no?

OpenStudy (anonymous):

uhm you can take an x out right?

hero (hero):

You have to find two numbers that multiply to get 3, but also add to get 4 \(m \times n = 3\) \(m + n = 4\) Can you guess the two numbers?

OpenStudy (anonymous):

1 and 3

OpenStudy (anonymous):

is the answer 0, 3, and 1?

hero (hero):

You're sort of jumping ahead. You shouldn't

hero (hero):

There's a reason we factor first before trying to find x

OpenStudy (anonymous):

Im just trying to find the answer real fast so i can see if i got it right in my review.

hero (hero):

You found the two numbers, so now we will replace 4 with 3 + 1 in the quadratic \(x^2 + 4x + 3\) to get: \(x^2 + (3 + 1)x + 3\) Then distribute the x: \(x^2 + 3x + x + 3\) Now notice that we can factor \(x\) from the first two terms, and 1 from the least two terms: \(x(x + 3) + 1(x + 3)\) And now (x + 3) is the remaining common factor so: \((x + 3)(x + 1)\)

hero (hero):

So now....that we put that back in place of the quadratic: \(0 = x(x + 3)(x + 1)\)

hero (hero):

Can you see what the values of x are now?

OpenStudy (anonymous):

so its just 3 and 1

hero (hero):

No, the next step is to take each term and set it equal to zero using the zero product property: x = 0 x + 3 = 0 x + 1 = 0 Obviously x = 0 is one of the solutions. Try solving for x for the other two.

OpenStudy (anonymous):

oh yeah duh, 0, -3, -1

hero (hero):

it's okay to be able to figure out (in the end) what the answer is, but you should try to learn how to do ALL the steps.